# Algebraic Topology Principles behind Topological Quantum Error Correction

### v1

@@ Appendix A Mathematical Foundation
A.1 Curve Homotopy and Transversality
@@
A.2 Simplexes and Cell Complexes
In section III.3, the concept of simplexes has been mentioned a few times, and the formal definition is presented here. We followed the definition from section 2.1 in [26]. An $n$-simplex can be considered as a generalized triangle consisting of $(n+1)$ points $p_{0},p_{1},\cdots,p_{n}$, that do not lie on the same hyperplane of dimension $m$ for all $m<n$. Also, if we normalize every point, they can be treated as vectors. Hence, a simplex can be represented as a collection of the length of the vectors $[p_{0},p_{1},\cdots,p_{n}]$, for instance, a standard $n$-simplex is
@@ A.2 Simplexes and Cell Complexes
Therefore, the chain complex of the cell complex $X$ can be written as $C^{cell}_{n}(X)=\langle C^{n}_{\sigma}\rangle=\langle A_{n}\rangle$. We name $X^{n}$ as the $n$-skeleton of the cell complex $X$, this definition is very useful when we discussed the TQEC codes on some higher-dimensional manifolds $X_{M}$ in Section IV.
A.3 Pairing and Cup Product
@@
A.4 The Betti Number and Homology
@@
A.5 Connected Sum Operation
In section III.3, when providing the examples, we mentioned that all 2-manifolds can be classified into two categories, where all orientable surfaces are homeomorphic to $gT^{2}$, and all non-orientable surfaces are homeomorphic to $gP^{2}$. And $gT^{2}:=(g-1)T^{2}\#T^{2}$ and $gP^{2}=(g-1)P^{2}\#\mathbb{R}P^{2}$, with the ##\## representing the connected sum operation. We would like to discuss the connected sum of two $n$-manifolds, $M$ and $N$.
@@ A.5 Connected Sum Operation
Appendix B TQEC codes on 2-dimensional cell complexes
As discussed in the concluding part of Section III.3 of the main text, we now seek to generalize the theorems presented therein and provide a framework for constructing TQEC codes on general two-dimensional cell complexes, which may not necessarily be embeddable in manifolds. To begin, we first recall a well-known theorem from algebraic topology, which serves as a foundational result for this generalization:
Theorem B.1.
[26, 27] For every Abelian group $G$ there is a 2-dimensional cell complex $X_{G}$ with $H_{1}(X_{G})\cong G$
This theorem is well established, and its proof can be found in standard algebraic topology textbooks, such as [26, 27], which also detail the construction of such cell complexes. It is important to note that in the theorem, $X_{G}$ is a two-dimensional cell complex that may not be embeddable in a 2-manifold due to its construction [26, 27]. While it is indeed possible to define TQEC codes on cell complexes that are not manifolds, the absence of manifold structure has significant implications. Specifically, for $X_{G}$ not embedded in a manifold, Poincaré duality II.1 does not hold. As a result, there is no direct correspondence between the cohomology group $H^{1}(X_{G};\mathbb{Z}_{2})$ and the set of loops on $X_{G}$. This lack of duality complicates the implementation of such codes, making them less straightforward compared to codes defined on 2-manifolds. We next cite another foundational theorem from cohomology theory [26, 27].
@@ Appendix B TQEC codes on 2-dimensional cell complexes
Each cohomology class in $H^{n}(X;F)$ can be uniquely identified with a class of linear functionals on the homology group $H_{n}(X;F)$, where each functional maps elements of $H_{n}(X;F)$ to $F$. Consequently, in principle, there exists a class of functionals capable of determining whether a loop $\gamma\in C_{1}(X;\mathbb{Z}_{2})$ satisfies $[\gamma]=0$ or $[\gamma]\neq 0$, which can be leveraged for TQEC purposes. However, in the absence of Poincaré duality, this class of linear functionals is not directly associated with loops in the dual cell complex $\bar{X}_{G}$. This contrasts with the case of 2-manifolds, where the dual cell complex can be embedded within the same manifold as a consequence of Poincaré duality. Therefore, according to this theorem, TQEC codes can, in theory, be constructed on more exotic two-dimensional cell complexes, albeit with increased implementation complexity.
@@
Appendix C Simulation on Klein Bottle Code for $\sigma_{X}$ Error
Figure 17: Logical versus physical $\sigma_{X}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different code distances assuming perfect syndrome measurements. (a) $d=\{8,14,20\}$ are even. (b) $d=\{7,13,19\}$ are odd. There is no statistically significant distinction between the performance of TQEC codes on $K$ or $T^{2}$ when $d$ is odd.
In this section, we present the simulation results for the logical $\sigma_{X}$ versus physical $\sigma_{X}$ error rates. As discussed in Section V.5 and illustrated in Figure 16(a), the shortest co-cycle $\beta\in C^{1}$ for both the Klein bottle and the torus remains identical for a given lattice dimension $L$, irrespective of whether $L$ is even or odd. Consequently, we predict that the logical $\sigma_{X}$ error rates for the toric and Klein bottle codes will be more or less the same under these conditions, as formerly discussed in Section V.5.
The methodology employed here is consistent with that described in Section V.2. Specifically, both the physical error rate and the corresponding logical error rate are collected and plotted for analysis. The physical error rate $p$ is sampled from a uniformly distributed random variable. For each code distance $d$, simulations are performed over 5000 runs for both the torus $T^{2}$ and the Klein bottle $K$. The error bars depicted in the plots represent 1-sigma uncertainties.

### v2

@@ Appendix B Mathematical Foundation
B.1 Curve Homotopy and Transversality
@@
B.2 Simplexes and Cell Complexes
In section III.3, the concept of simplexes has been mentioned a few times, and the formal definition is presented here. We followed the definition from section 2.1 in [26]. An $n$-simplex can be considered as a generalized triangle consisting of $(n+1)$ points $p_{0},p_{1},\cdots,p_{n}$, that do not lie on the same hyperplane of dimension $m$ for all $m<n$. Also, if we normalize every point, they can be treated as vectors. Hence, a simplex can be represented as a collection of the length of the vectors $p_{0},p_{1},\cdots,p_{n}]$, for instance, a standard $n$-simplex is
@@ B.2 Simplexes and Cell Complexes
Therefore, the chain complex of the cell complex $X$ can be written as $C^{cell}_{n}(X)=\langle C^{n}_{\sigma}\rangle=\langle A_{n}\rangle$. We name $X^{n}$ as the $n$-skeleton of the cell complex $X$; this definition is very useful when we discussed the TQEC codes on some higher-dimensional manifolds $X_{M}$ in Section IV.
B.3 Manifolds and Orientation
Before we present the definition of oriented manifolds and cell complexes, we first need to define charts and atlases on manifolds. These definitions are widely discussed in numerous mathematical textbooks, such as Ref. [34, 27].
Definition B.3 (Chart and (smooth) atlas).
Let $M$ be a topological space and fix $n\in\mathbb{N}$. A chart on $M$ is a pair $(U,\varphi)$ where $U\subset M$ is open and $\varphi:U\to\varphi(U)\subset\mathbb{R}^{n}$ is a homeomorphism onto an open subset.
A smooth atlas on $M$ is a collection $\mathcal{A}=\{(U_{\alpha},\varphi_{\alpha})\}$ of charts such that:
  • The domains cover $M$, i.e. $\bigcup_{\alpha}U_{\alpha}=M$.
  • Any two charts are smoothly compatible: for all $\alpha,\beta$ with $U_{\alpha}\cap U_{\beta}\neq\varnothing$, the transition map $\varphi_{\beta}\circ\varphi_{\alpha}^{-1}$ (defined on $\varphi_{\alpha}(U_{\alpha}\cap U_{\beta})$) is a smooth map between open subsets of $\mathbb{R}^{n}$.
Then, we can use the definition chart and atlas to define the orientability of a manifold.
Definition B.4 (Oriented smooth manifold).
Let $M$ be an $n$-dimensional smooth manifold. An orientation on $M$ is an equivalence class of atlases $\{(U_{\alpha},\varphi_{\alpha})\}$, where each $\varphi_{\alpha}:U_{\alpha}\to\mathbb{R}^{n}$ is a diffeomorphism onto its image, such that for any pair of overlapping charts $(U_{\alpha},\varphi_{\alpha})$, $(U_{\beta},\varphi_{\beta})$, the transition map $\varphi_{\beta}\circ\varphi_{\alpha}^{-1}:\varphi_{\alpha}(U_{\alpha}\cap U_{\beta})\to\varphi_{\beta}(U_{\alpha}\cap U_{\beta})$ has Jacobian determinant everywhere positive. The pair $(M,\mathcal{O})$, where $\mathcal{O}$ is such an orientation, is called an oriented smooth manifold.
Definition B.5 (Oriented cell complex).
Let $X$ be a cell complex and let $e^{n}$ be an $n$-cell attached via a characteristic map $\phi:D^{n}\to X$. An orientation of the cell $e^{n}$ is a choice of generator of the top relative homology group $H_{n}(D^{n},\partial D^{n};\mathbb{Z})\cong\mathbb{Z}$, or equivalently, a choice of orientation on the model ball $D^{n}$. An oriented cell complex is a cell complex equipped with a choice of orientation for every cell such that whenever $e^{n}$ is attached along $\phi|_{\partial D^{n}}:\partial D^{n}\to X^{(n-1)}$, the induced orientation on $\partial D^{n}$ agrees with the orientations of the (n1)(n-1)-cells it meets, in the sense that all attaching degrees are +1+1.
The definitions can be intuitively understood. A manifold is orientable if it admits a consistent global choice of local orientations, and non-orientable if no such global choice exists. Equivalently, non-orientability can be detected by the existence of a loop along which “clockwise” returns as “counterclockwise,” meaning local orientation flips after going around the loop, which is also equivalent to containing a subset homeomorphic to a Möbius strip.
Therefore, on Non-oriented manifolds, a global choice of local orientations is not well-defined, and the direction of intersections between 2 loops is also not well-defined. However, on oriented surfaces, since one can adopt a well-defined global choice of local orientations, the direction of intersection of several loops can be defined.
B.4 Pairing and Cup Product
@@
B.5 The Betti Number and Homology
@@
B.6 Connected Sum Operation
In section III.3, when providing the examples, we mentioned that all 2-manifolds can be classified into two categories, where all orientable surfaces are homeomorphic to $gT^{2}$, and all non-orientable surfaces are homeomorphic to $gP^{2}$. And $gT^{2}:=(g-1)T^{2}\#T^{2}$ and $gP^{2}=(g-1)P^{2}\#\mathbb{R}P^{2}$, with the #\# representing the connected sum operation. We would like to discuss the connected sum of two $n$-manifolds, $M$ and $N$.
@@ B.6 Connected Sum Operation
B.7 Universal Coefficient Theorem
We next cite another foundational theorem from cohomology theory [26, 27].
@@ B.7 Universal Coefficient Theorem
Each cohomology class in $H^{n}(X;F)$ can be uniquely identified with a class of linear functionals on the homology group $H_{n}(X;F)$, where each functional maps elements of $H_{n}(X;F)$ to $F$. Consequently, in principle, there exists a class of functionals capable of determining whether a loop $\gamma\in C_{1}(X;\mathbb{Z}_{2})$ satisfies $\gamma]=0$ or $\gamma]\neq 0$, which can be leveraged for TQEC purposes. However, in the absence of Poincaré duality, this class of linear functionals is not directly associated with loops in the dual cell complex $\bar{X}_{G}$. This contrasts with the case of 2-manifolds, where the dual cell complex can be embedded within the same manifold as a consequence of Poincaré duality. Therefore, according to this theorem, TQEC codes can, in theory, be constructed on more exotic two-dimensional cell complexes, albeit with increased implementation complexity.
@@
Appendix C Simulation on Klein Bottle Code for Pauli X Error
Figure 28: Logical versus physical $\sigma_{X}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different lattice dimensions assuming perfect syndrome measurements. (a) $L=\{8,14,20\}$ are even. (b) $L=\{7,13,19\}$ are odd. There is no statistically significant distinction between the performance of TQEC codes on $K$ or $T^{2}$ when $d$ is odd.
In this section, we present the simulation results for the logical $\sigma_{X}$ versus physical $\sigma_{X}$ error rates. As discussed in Section VII.5 and illustrated in Figure 27(a), the shortest co-cycle $\beta\in C^{1}$ for both the Klein bottle and the torus remains identical for a given lattice dimension $L$, irrespective of whether $L$ is even or odd. Consequently, we predict that the logical $\sigma_{X}$ error rates for the toric and Klein bottle codes will be more or less the same under these conditions, as formerly discussed in Section VII.5.
The methodology employed here is consistent with that described in Section VII.2. Specifically, both the physical error rate and the corresponding logical error rate are collected and plotted for analysis. The physical error rate $p$ is sampled from a uniformly distributed random variable. For each lattice dimension $L$, simulations are performed over 5000 runs for both the torus $T^{2}$ and the Klein bottle $K$. The error bars depicted in the plots represent 1-sigma uncertainties.